OpenStudy (anonymous):

What is the solution set for x squared -8x+4>-3? Walk me through how to solve this please? ( :

OpenStudy (anonymous):

so its x^2-8x+>-3?

OpenStudy (anonymous):

with a +4 after the 8

OpenStudy (anonymous):

sorry bout that so is it wanting to find the x though

OpenStudy (anonymous):

a solution set for x. My choices are 7<x<1 7<x>1 x is greater than 1 or x is greatr than 7 x <1 or x>7

OpenStudy (anonymous):

the first choice makes no sense

OpenStudy (anonymous):

neither does some of the math but how would you make x^2-8x+4 greater the a -3

OpenStudy (anonymous):

replace the x with a number that fit both the options and if it works there you go ex. 8^2-8(8)+4>3 i replaced it with an 8 which makes it 16-64+4>3 actually that would make it false cause 16-64=-48+4 = -44 so 3 would be greater

OpenStudy (anonymous):

thanks for the help but just guessed

OpenStudy (anonymous):

it was the second one lol heads up

OpenStudy (anonymous):

and btw youre very beautiful

OpenStudy (lyrae):

Another approach: \[x ^{2} - 8x + 4 > -3 \rightarrow x ^{2} - 8x + 7 > 0\]Assume \[x ^{2}−8x+7=0 \rightarrow x _{1} = 1, x _{2} = 7\] We know the basic shape of a possitive quadratic and the points where this specific quadratic intersects 0.From this it's quite easy to derive that the quadratic is grater than 0 when x < 1 or x > 7.